OpenStudy (anonymous):

LaRhonda is joining a gym next month. Gym A charges $40 for the membership fee plus the standard $25 per month, while Gym B has a membership fee of $20 in addition to the monthly charge of $35. How many months will LaRhonda need to attend the gym for the cost of the gyms to be the same? 10 months 6 months 4 months 2 months @gorv @Joel_the_boss

JoelTheBoss (joel_the_boss):

lol I really don't know xD

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@k_lynn

OpenStudy (anonymous):

@caitlinnr14

OpenStudy (k_lynn):

well to find the amount you need to pay for gym a would be the equation \[40+25x\] THen the amount you need to pay for gym B would be the equation \[20+35x\] Start plugging in numbers that you think would be logical for this problem and use it for both equations. For example, instead of x put 2 or something.

OpenStudy (anonymous):

I got it , the answer is 2, thanks

OpenStudy (k_lynn):

yep

OpenStudy (gorv):

let for x month

OpenStudy (gorv):

40+x*25 for A

OpenStudy (gorv):

that is cost for A

OpenStudy (gorv):

x month charge of B=x*35

OpenStudy (gorv):

total=20+x*35 for price to be same 20+x*35=40+x*25